# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering

### v1

@@ 3.3 Evaluation Metrics
In addition to the hallucination metrics, we have also included four metrics, following previous works (Edge et al., 2024; Wang et al., 2023) to quantify the desirable qualities for generating sensible responses. Given the complexities inherent in our open-ended QA setting, collecting and annotating ground-truth answers pose significant challenges. Therefore, we leverage LLMs as automatic reference-free evaluators to assess these metrics based on G-Eval (Liu et al., 2023), a framework of using LLMs with chain-of-thought (CoT) and a form-filling paradigm. To enhance G-Eval’s robustness and stability, we provide the relevant Wikipedia pages as context in the prompt when calculating the metrics. The metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context. Comprehensiveness: evaluates the extent to which the generated answer covers all aspects and details of the question. Correctness: measures how clearly and specifically the generated answer responds to the question. Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions. The detailed prompt for each metric can be found in Appendix A.1.
Models. We incorporate a range of widely used LLMs of different scales, including GPT-4o [^3] https://openai.com/index/hello-gpt-4o/, GPT-4o-mini [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024). We use the embedding model (text-embedding-3-small) from OpenAI for the G-retrieval process, considering cost and performance.
4 Exploring KG-augmented framework for reducing hallucination
Figure 2: Overview of KG-augmented framework for reducing hallucinations.
Overview. To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we investigated a range of methods and discovered that they can be encapsulated within a unified framework as shown in Figure 2. Our framework follows the paradigm of retrieval augmented generation (RAG) (Edge et al., 2024; Baek et al., 2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation. It has two components, i.e., Graph-guided retrieval (G-retrieval) and Graph-guided generation (G-generator), with a variety of algorithmic choices that combine to create diverse algorithms with different benefits for reducing hallucinations. We analyze our proposed strategies within each component in Section 5, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
Here we formalize the problem of context-relevant KG retrieval for open-ended question answering. Given a user query $q$, a generative pretrained language models (PLMs) first encodes the input tokens, and then models a conditional distribution $p(a|q)$ to generate an output response $a=[a_{1},\ldots,a_{T}]$. To explore whether KGs can help reduce hallucinations of PLMs, we introduce the retrieved knowledge from the KG, denoted $\mathcal{Z}\in\mathcal{G}$, which can take various forms (e.g., triplets, paths, subgraphs, as shown in Figure 2). This retrieved knowledge contains information relevant to the user query, and can either benefit the PLM at its knowledge boundaries (Baek et al., 2023b) or explicitly guide its reasoning process (Luo et al., 2023a; Sun et al., 2023). We explore which forms of knowledge should be extracted from KGs to best enhance the PLM. The distribution of the retrieved knowledge $\mathcal{Z}$ is modeled as $p(\mathcal{Z}|q,\mathcal{G})$, leading to a new likelihood for generating responses, expressed as $p(a|q)$.
where $p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G})$ represents the distribution of retrieving relevant knowledge elements from the KGs, and $p_{\phi}(a|q,\mathcal{Z}^{*})$ represents the distribution of leveraging the retrieved knowledge elements to generate the final responses, parameterized by $\phi$ and $\theta$, where $\mathcal{Z}^{*}$ is the optimal retrieved knowledge. Due to the potential exponential growth in the number of possible retrieved knowledge with graph size, efficient approximation methods are required. Consequently, the first line of Eq. 1 is approximated using the most probable retrieved knowledge $\mathcal{Z}^{*}$.
4.1 Graph-guided retrieval (G-retrieval)
The graph-guided retrieval phase focuses on extracting the most relevant elements (e.g., triplets, paths or subgraphs) from KGs to help answer the user query. Initially, the query is transformed into an embedding vector $\mathbf{q}\in\mathbb{R}^{d}$ using a language model encoder $f_{\text{LM}}(\cdot)$. For each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$, their respective embeddings $\mathbf{e}_{v}$ and $\mathbf{e}_{e}$ are computed using the same language model to map all the information into the same vector space, and their relevance to the query is quantified using cosine similarity as: $s_{v}=\frac{\mathbf{q}\cdot\mathbf{e}_{v}}{\|\mathbf{q}\|\|\mathbf{e}_{v}\|}$ and $s_{e}=\frac{\mathbf{q}\cdot\mathbf{e}_{e}}{\|\mathbf{q}\|\|\mathbf{e}_{e}\|}$. Next, we identify the most relevant nodes and edges for the query based on the corresponding similarity, to yield a set of ‘top-$k$ nodes/edges’.
The goal of G-retrieval is to retrieve the knowledge that encompasses as many relevant nodes and edges as possible, while keeping the size manageable. To this end, we leverage a ‘prize’ and ‘cost’ trade-off strategy to guide the retrieval process, as follows: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query. Specifically, we assign the top-$k$ nodes/edges with descending prizes values from $k$ to 1, while nodes and edges outside the top-$k$ receive a prize of 0. Formally: $p_{v}=\max(0,k-\text{rank}(v)+1)$ and $p_{e}=\max(0,k-\text{rank}(e)+1)$. (2) Cost allocation: to manage the retrieved knowledge size, we assign penalties as cost $C_{e}$ (by default assign to 1) during the expansion of the retrieved paths or subgraphs. These prizes and costs will be used in the retrieval process, by aiming to maximize the prizes while minimizing costs.
Path-retrieval: based on the defined prize assignment and cost allocation, the path-retrieval approach starts from a fix number of $k$ of high-prize nodes to construct sequences $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ by aiming to greedily maximize their scores: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$. We use a priority queue to iteratively extend paths by selecting the highest-scoring extensions while avoiding cycles and adhering to maximum path length and quantity constraints. Finally, the top-scoring paths are sorted and returned. The details of path-retrieval can be found in Appendix A.6.
Sub-graph retrieval: the sub-graph retrieval approach also leverages the prize assignment and cost allocation; we follow He et al. (2024) to optimize the process using Prize-Collecting Steiner Tree (PCST) algorithm. The PCST problem seeks to build a connected subgraph $\mathcal{S}$ that maximizes the total prizes of included nodes and edges while minimizing the total cost of the edges. The score of the subgraph is defined as $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{e\in E_{\mathcal{S}}}c_{e}$. Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
4.2 Graph-guided generation (G-generator)
After retrieving $\mathcal{Z}^{*}$, the graph-guided generation phase focuses on learning to leverage this knowledge to generate the corresponding output to answer the user query. The generation is modeled as a sequential decision-making process as follows, where at each time step $t$, the next token $a_{t}$ is generated by G-Generator, conditioned on the query $q$, the retrieved knowledge $\mathcal{Z}^{*}$, and all previously generated tokens $a_{0:t-1}$.
$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}\left(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}\right)$ (2)
5 Experiments
5.1 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
To explore whether KGs can help reduce hallucination in LLMs, we benchmark the LLMs in different settings. We use zero-shot and few-shot prompting as baselines without injecting external knowledge. In addition, compared to adding external knowledge from KGs, we also consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022) prompting as baselines. As for LLMs augmented with KGs, we consider three KG retrieval forms: triplets, paths, and subgraphs (as shown in Figure 2) for comparison, to study the impact of G-retrieval for reducing LLMs’ hallucinations. The results can be found in Table 2 and Figure 3. We obtain some intriguing findings:
@@ 5.1 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
GPT-4o - CoT+SC $82.90\%\pm 0.57\%$ $78.72\%\pm 0.61\%$ $73.64\%\pm 0.43\%$ $75.80\%\pm 0.75\%$ $89.12\%\pm 0.94\%$ $75.42\%\pm 1.31\%$
@@
Retrieving KG information can indeed mitigate factual errors in the responses. Methods integrating knowledge extracted from KGs show clear improvements in factual accuracy and comprehension scores compared to the baselines. For example, under Var-2 (triplet retrieval), GPT-4o achieves a FActScore of 72.55% $\pm$ 0.85%, which is a significant increase over the baseline score of 55.35% $\pm$0.95%. Moreover, these methods can be combined with strategies like CoT+SC, enhancing response quality with minimal increase in hallucination ratio. The radar chart in Figure 3 further emphasizes that in most query types, integrating knowledge retrieved from KGs mitigates the hallucination issue compared to baselines, particularly in query types such as “Evaluation and Reflection," “Outcome Prediction," and “Cause Explanation," which require more reasoning and analysis rather than merely listing information. The findings also apply to open-source models like mistral-7B-Instruct-v0.3 and Llama-3.1-8B-instruct, illustrating the consistency of reliability of integrating KGs in LLM reasoning.
@@
Directly performing reasoning in the LLM itself does not mitigate hallucinations. We benchmark the hallucination ratio of LLMs using internal reasoning strategies like Chain-of-thought and Self-consistency, as shown in Var-1 in Table 2. It shows that these methods can indeed improve response quality (i.e., correctness, comprehensiveness and empowerment) compared to baselines, but do not consistently improve factuality, and sometimes even diminish factuality. This showcases that relying solely on internal reasoning does not mitigate hallucination, emphasizing the importance of external knowledge to counter the hallucination issue.
Subgraph retrieval generally achieves best performance across different query types, especially for simpler queries. We demonstrate the performance of different retrieval methods across different query types in Figure 3, showing that subgraphs generally achieve best performance. Especially for simpler queries (“Character Description" and “Event Description" which generally don’t require intensive reasoning), subgraph retrieval methods achieves better performance compared to paths and triplets. Even for queries like “Relationship Explanation" and “Cause Explanation" which require step-wise reasoning, subgraph methods still demonstrate promising performance. This suggests that while different forms of retrieved knowledge each offer unique benefits suitable for specific types of queries, subgraphs provide consistently good performance broadly.
We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality. Specifically, we consider different perturbation methods discussed in Section 3.2 and control the perturbation level based on the percentage of KG edges being perturbed. We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 4.
@@ 5.2 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases. For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS is very high. Relation replacement has a relatively strong impact on ATS, i.e., on the semantics of the perturbed KGs. For SC2D and SD2 (which intuitively measure structural similarity), the four methods show similar levels of decrease in most cases.
@@
In conclusion, our study demonstrates that leveraging external knowledge from KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree. In real-world applications, KGs such as Wikidata and unlikely to experience severe perturbations due to their continuous updates and community-driven quality control, making our findings applicable in most real-world scenarios.
@@ 6 Conclusion
In this paper, we present OKGQA, a benchmark tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question answering scenarios. The benchmark extends the assessment from close-ended question answering to the open-ended setting, to support the assessment of hallucination in LLMs. To further mimic real-world scenarios where KGs may not have high quality, we propose another experiment setup OKGQA-P to assess model performance under conditions where KG’s semantics and structure are deliberately perturbed and contaminated. We conduct a series of experiments on OKGQA and OKGQA-P, analyzing the effectiveness of various retrieval methods and LLMs of different magnitudes, providing insights for further research and development. Our results underscore the significance of integrating KGs with LLMs to help reduce hallucination of LLMs, even in circumstances where the KGs are contaminated.
@@ Appendix A Appendix
A.1 Prompt List

### v2

@@ 4.1 Evaluation Metrics
[…]
In addition to the hallucination metrics, we also propose four metrics to quantify the desirable qualities for generating sensible responses (Edge et al., 2024; Wang et al., 2023). Given the complexities inherent in an open-ended setting, collecting and annotating ground-truth answers poses significant challenges. Therefore, we leverage LLMs as automatic reference-free evaluators to assess these metrics based on G-Eval (Liu et al., 2023), a framework of using LLMs with chain-of-thought (CoT) and a form-filling paradigm. To enhance G-Eval’s robustness and stability, we provide the relevant Wikipedia pages as context in the prompt when calculating the metrics. The metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context. Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question. Correctness: measures the clarity and specificity of the generated answer to the question. Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions. The detailed prompt can be found in Appendix C.
Models. We incorporate a range of widely used LLMs of different scales, including GPT-4o, GPT-4o-mini, Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024). We use the embedding model (text-embedding-3-small) from OpenAI for the G-retrieval process, considering cost and performance.
4.2 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
To explore whether KGs can help reduce hallucination in LLMs, we benchmark the LLMs in different settings. We use zero-shot and few-shot prompting as baselines without injecting external knowledge. In addition, compared to adding external knowledge from KGs, we also consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022) as baselines. For LLMs augmented with KGs, we consider three KG retrieval forms: triplets, paths, and subgraphs (as shown in Figure 3) to study the impact of G-retrieval for reducing LLMs’ hallucinations. The results are shown in Table 2 and Figure 4. We obtain some intriguing findings:
@@ 4.2 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
Retrieving KG information can indeed mitigate factual errors in the responses. Methods integrating knowledge extracted from KGs show clear improvements in factual accuracy and comprehension scores compared to the baselines. For example, under Var-2 (triplet retrieval), GPT-4o achieves a FActScore of 72.55% $\pm$ 0.85%, which is a significant increase over the baseline score of 55.35% $\pm$0.95%. Moreover, these methods can be combined with strategies like CoT+SC, enhancing response quality with minimal increase in hallucination ratio. The radar chart in Figure 4 further emphasizes that in most query types, integrating knowledge retrieved from KGs mitigates the hallucination issue compared to baselines, particularly in query types such as “Evaluation and Reflection,” “Outcome Prediction,” and “Cause Explanation,” which require more reasoning and analysis rather than merely listing information. The findings also apply to open-source models like mistral-7B-Instruct-v0.3 and Llama-3.1-8B-instruct, illustrating the consistency of the finding.
@@
Directly performing reasoning in the LLM itself does not mitigate hallucinations. We benchmark the hallucination ratio of LLMs using internal reasoning strategies like Chain-of-thought and Self-consistency, as shown in Var-1 in Table 2. It shows that these methods can improve response quality (i.e., correctness, comprehensiveness, and empowerment) compared to baselines, but do not consistently improve factuality, and sometimes even diminish factuality. This shows that relying solely on internal reasoning does not mitigate hallucination, emphasizing the importance of external knowledge to counter the hallucination issue.
@@
Subgraph retrieval generally achieves best performance across different query types, especially for simpler queries. We demonstrate the performance of different retrieval methods across different query types in Figure 4, showing that subgraphs achieve the best performance. Especially for simpler queries (“Character Description” and “Event Description” which do not require intensive reasoning), subgraph retrieval methods achieve better performance compared to paths and triplets. Even for queries like “Relationship Explanation” and “Cause Explanation” which require stepwise reasoning, subgraph methods still demonstrate promising performance. This suggests that while different forms of retrieved knowledge offer unique benefits for specific types of queries, subgraphs provide consistently good performance.
We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality. Specifically, we consider different perturbation methods discussed in §2.2 and control the perturbation level based on the percentage of KG edges being perturbed. We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 5. It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases. For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS remains higher compared to other settings. Relation replacement has a relatively stronger impact on ATS, i.e., on the semantics of the perturbed KGs. For SC2D and SD2 (which intuitively measure structural similarity), the four methods show similar levels of decrease in most cases.
@@ 4.3 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
Our findings on OKGQA-P illustrate that the effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance. We think that before this perturbation level at 50%, incorporating external knowledge from KGs can mitigate hallucinations in LLMs compared to baseline using CoT. In practical scenarios, platforms like Wikidata are less likely to be severely perturbed than 50% due to ongoing updates and community-based quality control, ensuring the relevance of our findings in real-world applications.
@@ 5 Conclusion
In this paper, we propose OKGQA and variant OKGQA-P, to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. Unlike existing benchmarks that focus primarily on closed ended tasks, OKGQA presents diverse open-ended question types that mirror the unpredictable nature of practical applications. We conduct a series of experiments and analyze the effectiveness of various retrieval methods and LLMs of different magnitudes, providing insights for further research. Our results underscore the significance of integrating KGs with LLMs to help reduce hallucination of LLMs, even in circumstances where the KGs are contaminated.
Limitations
Our proposed benchmark primarily use DBpedia as the knowledge source, which may not generalize well to testing scenarios requiring highly specialized or domain-specific knowledge. Testing domain-specific open-ended QA may require constructing sub-graphs from domain-specific KGs. In addition, the study assumes a static KG for reasoning and retrieval. In dynamic environments where knowledge is continuously updated, maintaining and integrating real-time changes remains a challenge and may requires further design.
@@ Appendix A Appendix
A.1 Intuition of Open-ended Questions?
In previous close-ended KGQA research Mihaylov et al. (2018); Yih et al. (2016); Talmor and Berant (2018), “hallucination” has not been considered at all. The primary reason is that close-ended KGQA models typically generate concise entity names as answers, which are directly compared against ground truth labels using metrics such as precision or accuracy. Any incorrect entity name is immediately reflected in the evaluation results, leaving little room for hallucination to manifest as a separate problem. While effective for simple tasks, we believe this closed-ended setting is too naive and fails to generalize to real-world, complex QA situations, where models must provide more detailed, nuanced responses Kantharaj et al. (2022).
In contrast, our work focuses on open-ended KGQA, where LLMs are prompted to generate more elaborate answers that include reasoning paths and supporting facts derived from the background KG. This broader output space not only provides a larger field to assess hallucination but also increases the likelihood of factual inaccuracies, as LLMs tend to produce more errors when generating longer and more complex responses. While hallucination is minimal in close-ended KGQA due to the short and factual nature of entity names, open-ended KGQA highlights the susceptibility of LLMs to hallucinate when generating rich, explanatory answers. By introducing this open-ended setting, OKGQA, we (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
A.2 Query Construction
In this section, we discuss the details of the query construction of OKGQA. It follows a template-based method with LLMs to generate a diverse range of queries. To ensure that the generated queries accurately represent real-world scenarios and complexities, we integrate the corresponding Wikipedia pages of each entity as contextual information for the LLMs’ generation. Furthermore, we follow a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 7. Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as an automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance. Then, these queries are manually evaluated and assigned a set of human-label scores $s_{\text{human}}$ corresponding to $s_{\text{auto}}$, normalized to the same range as $s_{\text{auto}}$. We then optimize the input instructions by iteratively generating instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$. This process quite mimics the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than labeling the output directly.
Figure 7: Human-in-the-loop of query construction.
We have three evaluators participating in the manual assessment of query quality. All of the evaluators are computer science majors with fluent English skills. As the evaluation centers on various linguistic metrics such as naturalness, relevance, specificity, novelty, and actionability, we only require the evaluators to possess a fundamental understanding of English without restricting their majors. We calculate the Pearson correlation coefficients between human and LLM scores as shown in Table 3. It shows that as the rounds progress, agreement between humans and LLMs increases, suggesting that iterative feedback improves alignment between human annotation and LLM responses.
Table 3: Pearson correlation coefficients between human and LLM scores across rounds.
Metric Round 1 Round 2 Round 3 Round 4
Naturalness 0.60 0.65 0.69 0.74
Relevance 0.55 0.59 0.64 0.70
Specificity 0.46 0.54 0.60 0.65
Novelty 0.49 0.57 0.63 0.67
Actionability 0.33 0.41 0.48 0.53
In addition, we also consider verifying the inter-rater reliability across three evaluators as shown in Table 4. We report the Cohen’s Kappa coefficient for each pair of evaluators as follows. Using the Landis and Koch (1977) interpretation guidelines, the Cohen’s Kappa coefficients for Naturalness and Relevance (ranging from 0.79 to 0.85) fall within the “Substantial” to “Almost Perfect” categories, indicating strong inter-rater reliability for these metrics. This reflects a shared understanding of the evaluation criteria, resulting in consistent ratings among evaluators. For Specificity, Novelty, and Actionability, the coefficients range from 0.58 to 0.68, placing them primarily in the “Moderate” to “Substantial” categories. These results suggest moderate reliability for these metrics, likely due to subjective interpretation and less clearly defined evaluation guidelines. Novelty, with lower coefficients around 0.61 to 0.63, highlights variability in ratings, suggesting that evaluators may have differing perspectives on what qualifies as novel (but the inter-rater reliability is still considered “Substantial”). Meanwhile, Actionability performs slightly better, nearing the “Substantial” range, indicating moderately consistent criteria.
Table 4: Cohen’s Kappa coefficient for various metrics.
Metric Evaluator 1 & 2 Evaluator 1 & 3 Evaluator 2 & 3
Naturalness 0.85 0.83 0.84
Relevance 0.81 0.79 0.80
Specificity 0.65 0.63 0.66
Novelty 0.60 0.58 0.61
Actionability 0.67 0.65 0.68
A.3 Personalized PageRank (PPR)
In this section, we discuss the details of the PPR algorithm used in §2.1 to prune the graph from DBPedia and concentrate on nodes most pertinent to the central nodes of interest. The PPR is calculated using the iterative formula:
$\mathbf{p}=\alpha\mathbf{A}^{\top}\mathbf{p}+(1-\alpha)\mathbf{s},$ (3)
[…]
A.4 Prize-Cost-based Path Retrieval
The path-retrieval method is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations. The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs. To efficiently manage the exploration of potential paths, we utilize a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first. The method starts by picking a number of starting nodes with high prizes. We then expand these starting points by exploring neighboring nodes. For each neighbor, the method calculates a new score. This score is the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost. If this neighbor hasn’t been visited before, which helps avoid looping, the algorithm adds this neighbor to the path. This new path is then added to the priority queue. This expansion keeps going until the path reaches a maximum length or can’t be extended further. The algorithm keeps track of paths already explored to avoid repetition and ensure paths don’t loop back on themselves. When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
A.5 KG Similarity Metrics
To measure how much the perturbed KGs are deviated from the original KG used in §2. We re-present the according metrics from (Raman et al., 2020) below:
Similarity in Degree Distribution (SD2)  SD2 also measures structural similarity between two KGs, while addressing SC2D’s ineffectiveness when the KGs’ entities have tiny local clustering coefficients (e.g., the item KG used by recommender systems is roughly bipartite). In such cases, SC2D is always close to one regardless of the perturbation method, thus rendering SC2D useless. Let $\mathbf{d}^{r}$ denote the $|\mathcal{E}|$-dimensional degree vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{d}^{r}$ is $\text{deg}(e_{i})$. Then, the mean degree vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{d}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{d}^{r}$ and $\mathbf{d}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{d}^{r}$, respectively. SD2 is defined as $f_{\text{SD2}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}}{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
Appendix B Related Work
Due to the stochastic decoding process of Large Language Models (LLMs), i.e., sampling the next token in the sequence, LLMs exhibit probabilistic behaviors: (1) potentially yielding varied outputs of the same input across different instances (Agrawal et al., 2023); (2) cannot accurately interpret phrases or terms when the context is vague and resides in a knowledge gap of the model. This will lead to outputs that may sound plausible but are often irrelevant or incorrect. This will lead to outputs that may sound plausible but are often irrelevant or incorrect. This “hallucinations” undermines the reliability of LLMs (Huang et al., 2023). One emerging research trend is enhancing LLMs through integrating external knowledge graphs (Agrawal et al., 2023). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023; Sui et al., 2022). Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent.
Researchers employ diverse strategies to augment the LLMs by integrating external KGs (Sui et al., 2024; He et al., 2024b). For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering. Wu et al. (2023) finds that converting these triples into textualized statements can further enhance LLM performance. StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs. Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al. (2022) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations. IR-CoT (Trivedi et al., 2022) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions. MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs. Similarly, RoG (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable reasoning in LLMs.
Appendix C Prompt List